OpenStudy (anonymous):

A tree cast a 25 foot shadow. At the same time of day, a 6 foot man standing near the tree cast a 9 foot shadow. what is the approximate height of the tree to the nearest foot ?

OpenStudy (misty1212):



OpenStudy (misty1212):

use ratios maybe \(\frac{6}{9}=\frac{x}{25}\)

OpenStudy (danjs):



OpenStudy (danjs):

similar triangles

OpenStudy (danjs):

the sides are proportional, similar because they both have Angle-Angle - Angle (3 same angles)

OpenStudy (danjs):

So 9 to 25 is the same as 6 to h

OpenStudy (danjs):

\[\frac{ 9 }{ 25 } = \frac{ 6 }{ h }\] 9h = 25*6 = 150 h=150/9 = about 16.67 ft

OpenStudy (anonymous):

How do i find the nearest foot ... the options are a)16 feet b)17 feet c)18 feet d)19 feet

OpenStudy (danjs):

17

OpenStudy (anonymous):

Oh okay thank you I get it

OpenStudy (danjs):

always draw pictures and label everything you are given, makes it much easier to think through

OpenStudy (anonymous):

Thank you for helping me, it greatly appreciated. :)

OpenStudy (danjs):

no problem, anytime